Please can you help me in this question because am having alot of problem...
there it says complete the table of values for y=x-2-4x

Answer by praseenakos@yahoo.com(507)

Question:

Please can you help me in this question because am having alot of problem...
there it says complete the table of values for y=x-2-4x

Answer:

As per the given question, you have to find the values of y corresponding to some values of x.

here you have given that, y = x - 2 - 4x

You can simplify this expression like this...

That is y = -2 -5x

Now give some values for x. (If you are given some values for x with the question, then you can use those values....otherwise take some arbitrary values...)

When x = 0, y = -2 - 5*0 = -2 - 0 = 0
When x = 1, y = -2 - 5*1 = -2 - 5 = -7
When x = 2, y = -2 - 5*2 = -2 - 10= -12
when x = 3, y = -2 - 5*3 = -2 - 15 = -17


Like wise you can find out the values of y corresponding to the valuies of x.


Hope you found the explanation useful.

Regards.

Praseena.